answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is a event occurs last in the Apology?

The Apology is an interpretation of the philosopher Socrates' speech by fellow philosopher Plato. At the end of the speech, Socrates dies.

Probability of rolling a 3 twice with 2 rolls on a cube?

Probability of a 3 on the first roll = 1/6

Probability of a 3 on the second roll = 1/6

Probability of both of them = (1/6) x (1/6) = 1/36 = 0.0277777 = 2.78% (rounded)

Event A has probability 0.5 event B has probability 0.6 which of the following must be true?

the probabilty of both events is true. but which is most reliable is probabilty of B as it is more near to 1( total probabilty of any event)

A whole number consists of the digits 2 4 6 8 and 0 If each digit is used only once how many whole numbers are possible?

There are 5 digits in 24680; think of them as 5 compartments. You can select 5 numbers for the 1st compartment, 4 numbers for the 2nd compartment, 3 numbers for the 3rd compartment, 2 numbers for the 4th compartment, and 1 number for the 5th compartment. So, the answer is 5*4*3*2*1 or 5! or 120

What are the odds of rolling five ones with a dice in six tries?

You can do it in 6 ways:

11111x

1111x1

111x1x

11x111

1x1111

x11111

Each has the same probability: (1/6)5 x (5/6) = 1/6 x 1/6 x 1/6 x 1/6 x 1/6 x 5/6 = 5/46656

So the probability is 5/46656 x 6 = 5/7776

You can use the binomial theorem so answer more complex questions like these.

What is the number of distinguishable permutations of the group of letters in the word VILLAIN?

Permutations of 7 letters = (7 x 6 x 5 x 4 x 3 x 2 x 1)

Divide by 2 because swapping the two 'I's is indistinguishable.

Divide by 2 because swapping the two 'L's is indistinguishable.

Distinguishable permutations = (7 x 6 x 5 x 4 x 3 x 2) / 4 = 1,260

What is the probability of obtaining atleast one tail when a coin is flipped six times?

you are looking for the probability of getting one tail, two,......., and six

this equivalent to saying 1 - the probability of not getting any tails (or getting 6 heads = (1/2)^6).

P(X>=1) = 1 - (1/2)^6=

Three cards are chosen at random from a standard deck of cards without replacement What is the probability of getting 3 aces?

The probability of drawing the first ace is 4 in 52. The probability of getting the second ace is 3 in 51. The probability of getting the third ace is 2 in 50. The probability, then, of drawing three aces is (4 in 52) times (3 in 51) times (2 in 50), which is 24 in 132600, or 1 in 5525, or about 0.0001810

What is the probability of getting a sum less than 6 if you roll two number cubes?

Out of a total of 36 combinations (1-1, 1-2, 1-3, ... 2-1, 2-2, ... 6-6), you have the following favorable possibilities:

1-1

1-2

1-3

1-4

2-1

2-2

2-3

3-1

3-2

4-1

The total is 10 combinations out of 36, so the probability is 10/36.

What is the probability of rolling an even number on a die?

Total numbers on a cube = 6

Even numbers on a cube = 3

Probability of rolling an even number on one fair cube = 3/6 = 50% .

If you roll a number cube twice what is the probability of rolling a 6 and a 2?

Assuming that a dice roll is purely random, there is a 1 in 6 probability of landing on any number. Since the second roll depends on the probability of the first, they factor together:

First Roll: 1/6

Second Roll: (1/6)(1/6) = 1/36

If you were rolling both dice at once, however, the math would be completely different.

Probability of getting an even product when rolled 2 dice?

I'm assuming you multiply the numbers rolled. Thus the rules of even and odd numbers come into play. The only way to get an odd product is to roll two odd numbers. The probability desired is 1 minus the probability of rolling two odd numbers.

The odds of rolling an odd number is 1/2. Since the rolls are independent events, the odds of rolling two odd numbers is (1/2)2 = 1/4

So the probability of getting an even product is:

1 - 1/4 = 3/4

There are 4 red 4 black and 4 white marbles in a box how many do you have to take out without looking to be sure you have 2marbles of the same color?

4:

Because there will be at least 1 of two kinds of marbles, and 2 of the other.

Ex: 1 black, 1 red, and 2 whites.

What effect does sample size have on the estimate of the mean?

the larger you r sample size the better your estimate. imagine take the height of person to estimate the average high of an adult male. would one person's height be a good estimate, or would taking the average height of 100, or 5000 adult males will produce a better estimate?

Theoretical probability of rolling a factor of 10 on a number cube?

The factors of 10 are the numbers that divide 10 evenly: 1, 2, 5 and 10. To answer your question, you have to figure out what the probability of rolling one of these numbers is on a number cube.

If IQ scores are normally distributed with a mean of 100 and a standard deviation of 17 and if the population of the world is 6575000000 how many geniuses are there in the world today?

I'm assuming "genius" means an unusual value (more than two standard deviations) in the right tail. This is (100 - 95.45)/2 = 2.275 percent, approximately. So:

6,575,000,000 * 0.02275 = 149,581,250 geniuses ... not that special, huh?

Trending Questions
What is represenatative sample of population? What is EIA? Number that describes a certain event will occur? What is the probability of rolling a number greater than three if you roll a die two times? Why is the penny brown? Why belong exponential family for poisson distribution or geometric distribution? What is the probability of pulling a king of hearts out of a deck? What are the odds of rolling five ones with a dice in six tries? What can be said about a set of data when its standard deviation is small but not zero? If you toss a coin 8 times there are 256 possible outcomes so how many of these outcomes have exactly 3 heads? How is using survey one way to determine experimental probability? Why is it that only one normal distribution table is needed to find any probability under the normal curve? Flip coin 50 times and toss number cube 5o times what is the probability.? When You roll 1 die Find Pthe number rolled is odd? The probability that a given student will show up for the test is 0.9 if there are 27 students in the class what is the probability that everyone will show up? What is the study of probability called? Why were you born a girl? What is the probability of exactly 3 girls out of three children? If a sample of 4 different calculators is randomly selected from a group containing 11 that r defective n 30 that have no defects what is the probabilty that atleast 1 of the calculators is defective? What is the probability of drawing 2 similar cards from a deck of 5?